I, The reaction A→B+C has an activaれon energy CEA OF $5.3.Draus reachion londscape LA vs Rx n Coordl nate 2. n the previous problem how mueh would the Rate ncrease e the temp was raiseel trom 27 c o S3.4c ExpressYour ansuer as a statement
Answer #1

1) The reaction has a positive energy of activation. This means that the transition state has a higher energy than the reactants and the products. The free energy diagram is drawn below.

2) The Arrhenius equation is written as

ln k2/k1 = -Ea/R*(1/T2 – 1/T1)

where k1 and k2 are the reaction rates at temperatures T1 and T2.

Given T1 = 27°C = (27 + 273) K = 300 K and T2 = 53.4°C = (53.4 + 273) K = 326.4 K, plug in values and get

ln k2/k1 = -(55.3 kJ/mol)/(8.314 J/mol.K)*[1/(326.4 K) – 1/(300 K)]

= -[(55.3 kJ/mol)*(1000 J)/(1 k)]/(8.314 J/mol.K)*(-0.0002696 K-1)

= 1.7932

=====> k2/k1 = exp (1.7932)

=====> k2/k1 = 6.008 ≈ 6.0

=====> k2 = 6.0*k1

The rate of the reaction increases 6 times (ans).